PLEASE Help!
I have to answer these
An optical instrument is required to observe stars beyond the sixth magnitude, the limit of ordinary vision, however, even optical instruments have their limitations. The limiting magnitude L of an optical telescope with lens diameter D (in inches) is given by
L=8.8+5.1logD
a) Find the limiting magnitude for a homemade 6-inch reflecting telescope
Thank you so much!!

Answer by ankor@dixie-net.com(13486)

An optical instrument is required to observe stars beyond the sixth magnitude, the limit of ordinary vision, however, even optical instruments have their limitations. The limiting magnitude L of an optical telescope with lens diameter D (in inches) is given by
L = 8.8 + 5.1logD
:
a) Find the limiting magnitude for a homemade 6-inch reflecting telescope.
:
Just replace D with 6 and solve it
L = 8.8 + 5.1*log(6)
L = 8.8 + 5.1 * .778
L = 8.8 + 3.968857
L = 12.8 limiting magnitude